OpenStudy (anonymous):

Which of the following have a greater rate of change than the equation y = x + 3?

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the real steep line has a larger slope the one in the picture

OpenStudy (anonymous):

is that one of them @misty1212

OpenStudy (misty1212):

yes the picture B can you pick more than one, or only one?

OpenStudy (misty1212):

the slope of C is \[\frac{11-2}{8-2}=\frac{9}{6}=\frac{3}{2}\] that is also bigger than one

OpenStudy (anonymous):

more than one @misty1212